5 6 7 8 9 10 10.95 12 14.4 17.28 20.74 24.88 1

Greek Fonts for Textual Use

B Hamilton Kelly^{∗} Royal Military College of Science Shrivenham, SWINDON, SN6 8LA United Kingdom

December 15, 2004

Needing to write some Greek examples for my Adult Education evening class, and being ignorant (at the time) of the existence of Silvio Levy’s excellent work, I created a selection of driver files which used the definitions of the maths font Greek characters (assembled into GRKTXT.MF), together with some other useful bits. I used these to generate a Greek text font in normal and bold weights, and also in italic and typewriter styles.

The characters are mapped to the intuitive Roman alphabet equivalents, as shown in Table ??, with the use of some ligatures to handle Greek letters traditionally represented by two Roman letters. These mappings are not everywhere the same as Silvio Levy’s; in particular, note that he uses ‘H’ to represent ‘Y’^{1} (which is intuitive as a capital, but not in lower-case [‘h’≄‘y’]) — I feel that my transliteration here makes the “English” Greek more readable, since most of it tends to be in lower-case. (Incidentally, I originally started with more ligatures, for example, ‘ks’ and ‘rh’ for ‘x’ and ‘r’ respectively; the latter caused a minor disaster with the Greek for ‘I start’ (arhízw) which has to be input as

‘arh\'izw’!)

Since I was interested only in Modern (Demotic) Greek, which adopted the “one-accent” system in 1982, I made no special provision for accents: in particular, there is no simple way of generating breathing marks. I just use the normal L^{A}TEX \' before the accented letter: of course, this puts the accent on capital letters in the wrong place (Ý whereas it should be ´ Y), but one doesn’t always need to accent capitals, and \'{} works when necessary. I recognize that using this form of accenting precludes any attempt by TEX to hyphenate the Greek words, but I’m pretty clueless on where to hyphenate Greek in general! Some time, I ought to get around to making ‘;’ into a sentence-ending punctuation

∗Because I realize that “double-barrelled” names cause confusion in the U.S., and since mine is even more unusual in that it does not contain a hyphen, I should point out that my surname is ‘Hamilton Kelly’, and not just ‘Kelly’! (First name is ‘Brian’, by the way.)

1The \grbf font has been used in all the in-text examples in this paper

Greek Characters  Greek Characters
Letter  Typed  Letter  Typed

                A  A  a  a
                B  B  b  b
                G  G  g  g
                D  D  d  d
                E  E  e  e
                Z  Z  z  z
                Y  Y  y  y
               TH  TH or Th  th  th
                I  I  i  i
                K  K  k  k
                L  L  l  l
               M  M  m  m
                N  N  n  n
                X  X  x  x
                O  O  o  o
                P  P  p  p
                R  R  r  r
                S  S  s  s

                                              \s

                T  T  t  t
                U  U  u  u
               PH  PH or Ph  ph  ph
                H  H  h  h
               PS  PS or Ps  ps  ps
               W  W  w  w

Table 1: Transliteration from Keyboard Input to Greek

mark (since Greek uses it where English uses ‘?’), and perhaps provide a raised dot for the Greek semi-colon; at present, all punctuation marks are in the conventional (Roman) positions, without transliteration, so you type a ‘;’ to get the Greek question mark.

To facilitate the use of these fonts, I have designed a L^{A}TEX style file option GREEK.STY, which ought to be usable with any normal L^{A}TEX style; at present, it is not suitable for SliTEX. The commands \gr, \grbf, \grit and \grtt are defined, and will automatically track any L^{A}TEX size changing command (i.e. \Large will invoke a 14pt \rm font in a 10pt document style, and then selecting \grbf will select the corresponding 14pt Greek bold font). Whenever a Greek font definition is in force, the macro \s is used to generate the variant of lowercase \sigma required at the ends of words. (Silvio Levy’s font takes care of all this automatically, however, I’m pretty sure that my LN03 printer driver would fail if I needed more than 188 characters from his Greek fonts, because each TEX font needs to be mapped to a single LN03 internal font, which has a maximum capacity of two 94-character “chunks” in the GL and GR terminology of ANSI X-3.64.)

In the last few days I’ve been getting my hands dirtier with , and have changed the programs for a number of the characters; this was necessary because they still exhibited that they were originally “tuned” for maths setting. The characters changed are as follows:

• The lower-case \alpha was too wide and also extended right up to the right-hand edge of the bounding box; this has been shrunk by some judicious tweaking of points defining the pen stroke.

• The characters \iota and \mu also came right up to their right-hand sides; they have simply been redefined in a box one u wider.

• The character \nu had a vertical left stroke; this has been curved and the serif (if any) blended in.

• The character \lambda was too wide: fine for denoting wavelength and the like, but not very pretty in text; its width has been shrunk by one u.

• The character \kappa is fine for maths, but not the form ordinarily found in Greek text; what is wanted is something like the \varkappa provided by AMSTEX. My first thought was to pirate that character from YSYMBOL.MF (which we have) but of course that’s for the old MF-in-SAIL; I have written something in 84 which approximates that character (please don’t look too closely at the awful code!)

There is considerable scope for further work:

• As mentioned above, because the Greek character definitions were stolen from the maths fonts, spacing of some other characters is not yet perfect; furthermore, the actual form of some other characters (notably z and x) is not that normally expected in Greek text. Eventually, I shall have to tweek them myself, or steal the appropriate bits of Levy’s fonts. (As mentioned above, I feel that use of his full 256-character font would cause problems at this site.)

• To make GREEK.STY work with SliTEX as well, I shall have to re-write SFONTS.TEX to use a font size changing strategy akin to that used by LFONTS.TEX, this will also permit the use of CYRILLIC.STY (which we have also modified to track automatically L^{A}TEX size changes) with SliTEX; quite frankly, I could never understand why SFONTS.TEX preloads all those fonts, when load-on-demand would seem much more useful for the contexts in which SliTEX is used.

• I haven’t made any “guillemets” style quotation marks.

Finally, an example; I don’t know whether Silvio Levy asked the copyright holder’s permission to reproduce his example from Kazantzakis’ “Symposium”, but I haven’t! I have, however, followed modern usage of the one-accent system. (I have also corrected what I presume to be a “typo”; it may be, of course, that Kazantzakis actually wrote what was printed in Levy’s article, but I feel that m’ ézwsan [“encircled me”] makes more sense when referring to lightning than m’ éxwsan [“evicted me”] — my Megálo epítomo lexikón ty Ellyniký glẃssa [ca. 1500pp.] doesn’t suggest any other sensible meaning for this verb!)

Ty stigmý toúty niẃthw póso barú ’nai to mustýrio ty xomológysy. W tẃra, kaneí den xérei pẃ pérasa ta duó hrónia mou sto ´ Agion ´ Oro. Oi phíloi mou tharroún pw pýga na dw buzantiná konísmata ý apó mustikopátheia na zýsw miá perasmény epohý. Kai tẃra, na, ntrépomai na milýsw.

Pẃ na to pw; Thumoúmai éna anoixiátiko deilinó, pou katébaina ton Taúgeto, miá xaphniký thúella me kúklwse kontá stou Pentauloú. Tóso phoberó anemosíphouna, pou épesa katagý giá na mýn gkremistẃ. Oi astrapé m’ ézwsan oloúthe ki ékleisa ta mátia mýn tuphlwthẃ, kai katáhama, pístoma, perímena. ´ Olo to panúpsylo bounó étreme, kai duó élata dípla mou tsakístykan ap’ ty mésy kai bróntyxan hámou. ´ Eniwtha to theiáphi tou keraunoú ston aéra, kai xaphniká xéspase y mpóra, épesen o ánemo, kai hontré, thermé stále brohý htúpysan ta dentrá kai to hẃma. To thumári, y throúmpa, to phaskómylo, to phliskoúni, htupyména ap’ to neró, tínaxan ti murwdié tou ki óly y gy múrise.

And here is the text input to generate that:

\begin{flushleft}\parskip 5pt plus 1pt
{\gr Ty stigm\'y to\'uty ni\'wthw p\'oso bar\'u
'nai to must\'yrio ty\s\ xomol\'ogysy\s.  W\s\ t\'wra,
kane\'i\s\ den x\'erei p\'w\s\ p\'erasa ta du\'o hr\'onia mou
sto \'{}Agion \'{}Oro\s.  Oi ph\'iloi mou tharro\'un pw\s\
p\'yga na dw buzantin\'a kon\'ismata \'y ap\'o
mustikop\'atheia na z\'ysw mi\'a perasm\'eny epoh\'y.  Kai
t\'wra, na, ntr\'epomai na mil\'ysw.
P\'w\s\ na to pw\@;  Thumo\'umai \'ena anoixi\'atiko
deilin\'o, pou kat\'ebaina ton Ta\'ugeto, mi\'a xaphnik\'y
th\'uella me k\'uklwse kont\'a stou\s\ Pentaulo\'u\s.
T\'oso phober\'o\s\ anemos\'iphouna\s, pou \'epesa
katag\'y\s\ gi\'a na m\'yn gkremist\'w.  Oi astrap\'e\s\
m'~\'ezwsan olo\'uthe ki \'ekleisa ta m\'atia m\'yn
tuphlwth\'w, kai kat\'ahama, p\'istoma, per\'imena.  \'{}Olo

to pan\'upsylo boun\'o \'etreme, kai du\'o \'elata d\'ipla
mou tsak\'istykan ap' ty m\'esy kai br\'ontyxan h\'amou.
\'{}Eniwtha to thei\'aphi tou kerauno\'u ston a\'era, kai
xaphnik\'a x\'espase y mp\'ora, \'epesen o \'anemo\s, kai
hontr\'e\s, therm\'e\s\ st\'ale\s\ broh\'y ht\'upysan ta
dentr\'a kai to h\'wma.  To thum\'ari, y thro\'umpa, to
phask\'omylo, to phlisko\'uni, htupym\'ena ap' to ner\'o,
t\'inaxan ti\s\ murwdi\'e\s\ tou\s\ ki \'oly y gy\s\
m\'urise.}
\end{flushleft}

(Note particularly the ‘\@;’ to make the Greek question mark end the sentence; also the use of ‘\'{}’ to put the accents before the capital letters.)

Here it is again, but this time in the italic face:

Ty stigmý toúty niẃthw póso barú ’nai to mustýrio ty xomológysy. W tẃra, kaneí den xérei pẃ pérasa ta duó hrónia mou sto ´ Agion ´ Oro. Oi phíloi mou tharroún pw pýga na dw buzantiná konísmata ý apó mustikopátheia na zýsw miá perasmény epohý. Kai tẃra, na, ntrépomai na milýsw.

Pẃ na to pw; Thumoúmai éna anoixiátiko deilinó, pou katébaina ton Taúgeto, miá xaphniký thúella me kúklwse kontá stou Pentauloú. Tóso phoberó anemosíphouna, pou épesa katagý giá na mýn gkremistẃ. Oi astrapé m’ ézwsan oloúthe ki ékleisa ta mátia mýn tuphlwthẃ, kai katáhama, pístoma, perímena. ´ Olo to panúpsylo bounó étreme, kai duó élata dípla mou tsakístykan ap’ ty mésy kai bróntyxan hámou. ´ Eniwtha to theiáphi tou keraunoú ston aéra, kai xaphniká xéspase y mpóra, épesen o ánemo, kai hontré, thermé stále brohý htúpysan ta dentrá kai to hẃma. To thumári, y throúmpa, to phaskómylo, to phliskoúni, htupyména ap’ to neró, tínaxan ti murwdié tou ki óly y gy múrise.

And again, but in the bold extended form:

Ty stigmý toúty niẃthw póso barú ’nai to mustýrio ty xomológysy. W tẃra, kaneí den xérei pẃ pérasa ta duó hrónia mou sto ´ Agion ´ Oro. Oi phíloi mou tharroún pw pýga na dw buzantiná konísmata ý apó mustikopátheia na zýsw miá perasmény epohý. Kai tẃra, na, ntrépomai na milýsw.

Pẃ na to pw; Thumoúmai éna anoixiátiko deilinó, pou katébaina ton Taúgeto, miá xaphniký thúella me kúklwse kontá stou Pentauloú. Tóso phoberó anemosíphouna, pou épesa katagý giá na mýn gkremistẃ. Oi astrapé m’ ézwsan oloúthe ki ékleisa ta mátia mýn tuphlwthẃ, kai katáhama, pístoma, perímena. ´ Olo to panúpsylo bounó étreme, kai duó élata dípla mou tsakístykan ap’ ty mésy kai bróntyxan hámou. ´ Eniwtha to theiáphi tou keraunoú ston aéra, kai xaphniká xéspase y mpóra, épesen o ánemo, kai hontré, thermé stále brohý htúpysan ta dentrá kai to hẃma. To thumári, y throúmpa, to phaskómylo, to phliskoúni, htupyména ap’ to neró, tínaxan ti murwdié tou ki óly y gy múrise.

And one last time, in the Greek typewriter font:

Ty stigmý toúty niẃthw póso barú ’nai to mustýrio ty xomológysy. W tẃra, kaneí den xérei pẃ pérasa ta duó hrónia mou sto ´ Agion ´ Oro. Oi phíloi mou tharroún pw pýga na dw buzantiná konísmata ý apó mustikopátheia na zýsw miá perasmény epohý. Kai tẃra, na, ntrépomai na milýsw.

Pẃ na to pw; Thumoúmai éna anoixiátiko deilinó, pou katébaina ton Taúgeto, miá xaphniký thúella me kúklwse kontá stou Pentauloú. Tóso phoberó anemosíphouna, pou épesa katagý giá na mýn gkremistẃ. Oi astrapé m’ ézwsan oloúthe ki ékleisa ta mátia mýn tuphlwthẃ, kai katáhama, pístoma, perímena. ´ Olo to panúpsylo bounó étreme, kai duó élata dípla mou tsakístykan ap’ ty mésy kai bróntyxan hámou. ´ Eniwtha to theiáphi tou keraunoú ston aéra, kai xaphniká xéspase y mpóra, épesen o ánemo, kai hontré, thermé stále brohý htúpysan ta dentrá kai to hẃma. To thumári, y throúmpa, to phaskómylo, to phliskoúni, htupyména ap’ to neró, tínaxan ti murwdié tou ki óly y gy múrise.